# The FinBen: An Holistic Financial Benchmark for Large Language Models

### v1

@@ 2.2 Spectrum II: Advanced Cognitive Engagement
Generation. The generation task gauges the models’ proficiency in producing coherent, informative, and relevant text outputs, involving the Crystallized Intelligence (Gc). We focus on the text summarization task utilizing the ECTSUM Mukherjee et al. (2022) dataset for summarizing earnings call transcripts and the EDTSUM Zhou et al. (2021) dataset for abstracting financial news articles into concise summaries. It’s evaluated using ROUGE scores Lin (2004), BERTScore Zhang et al. (2019), and BART Score Yuan et al. (2021), metrics that quantify to measure the alignment, factual consistency, and information retention between machine-generated and expert summaries.
@@
Forecasting. The forecasting task leverages Fluid Intelligence (Gf), challenging models to adaptively predict future market and investor behaviors from emerging patterns. It includes 12 datasets from 5 forecasting tasks. 1) Stock movement prediction focuses on forecasting stock directions as either positive or negative, based on historical prices and tweets, utilizing three datasets: BigData22 Soun et al. (2022), ACL18 Xu and Cohen (2018) and CIKM18 Wu et al. (2018). 2) Credit scoring classifies individuals as "good" or "bad" credit risks using historical customer data, employing datasets including: German Hofmann (1994), Australia Quinlan and LendingClub Feng et al. (2023). 3) Fraud detection involve categorizes transactions as "fraudulent" or "non-fraudulent", using two datasets: ccf Feng et al. (2023) and ccFraud Feng et al. (2023). 4) Financial distress identification aims to predict a company’s bankruptcy risk, using the polish Feng et al. (2023) and taiwan dataset Feng et al. (2023). 5) Claim analysis anonymizes client data for privacy, labeling a "target" to indicate claim status, using two datasets: PortoSeguro Feng et al. (2023) and travelinsurance Feng et al. (2023). F1 score and Matthews correlation coefficient (MCC) Chicco and Jurman (2020) are used for evaluating these tasks.
2.3 Spectrum III: General Intelligence
@@
Trading. Strategic decision-making in Trading Punt (2017), categorized under Spectrum III, is the pinnacle task for financial LLMs, emphasizing their use of General Intelligence (g). This task evaluates the model’s proficiency in synthesizing diverse information to formulate and implement trading strategies, a challenge even for experts, representing the highest level of cognitive capability in financial analysis. The SOTA financial LLM agent FinMem Yu et al. (2023) are used to evaluate LLMs on sophisticated stock decisions, based on the FinTrade dataset we curated of seven major stocks, simulating real-world trading through historical prices, news, and sentiment analysis. Performance is measured by Cumulative Return (CR) Ariel (1987), Sharpe Ratio (SR) Sharpe (1998), Daily (DV) and Annualized volatility (AV) Zhou et al. (2023), and Maximum Drawdown (MD) Magdon-Ismail and Atiya (2004), offering a comprehensive assessment of profitability, risk management, and decision-making prowess.
@@ 3 Evaluation
We evaluate the zero-shot and few-shot performance of 15 representative general LLMs and financial LLMs on the FinBen benchmark, including: 1) ChatGPT: An instruction-following LLM with 175B parameters developed by OpenAI. 2) GPT-4 OpenAI (2023b): A powerful instruction-following LLM with approximately 1T parameters, proposed by OpenAI. 3) Gemini Pro Team et al. (2023): A multimodal AI LLM with 50T parameters, released by Google. 4) LLaMA2-70B Touvron et al. (2023): An instruction-following LLM with 70B parameters developed by MetaAI. 5) ChatGLM3-6B Du et al. (2022): A conversational LLM with 6B parameters, jointly released by Zhipu AI and Tsinghua KEG. 6) Baichuan2-6B Baichuan (2023): An open-source LLM with 6B parameters, launched by Baichuan Intelligent Technology. 7) InternLM-7B Team (2023b): An open-sourced 7B parameter base model tailored for practical scenarios, proposed by SenseTime. 9) Falcon-7B Almazrouei et al. (2023): A 7B parameter causal decoder-only LLM model trained on 1500B tokens of RefinedWeb enhanced with curated corpora. 10) Mixtral 8$\times$7B Jiang et al. (2024): A LLM with the Sparse Mixture of Experts (SMoE) architecture. 11) Code Llama-7B Roziere et al. (2023): An open-source LLM model for generating programming code, launched by Meta AI with 7B parameters. 12) FinGPT Yang et al. (2023a): An 7B instruction finetuned financial LLM with sentiment analysis tasks. 13) FinMA-7B Xie et al. (2023b): An 7B instruction finetuned financial LLM with multiple NLP and forecasting tasks. 14) DISC-FinLLM Chen et al. (2023c): An open-sourced financial LLM, fine-tuned from Baichuan-13B-Chat Baichuan (2023). 15) CFGPT Li et al. (2023a): An open-source LLM, specifically designed for the financial sector and trained on Chinese financial datasets, which comprises 7 billion parameters. All experiments are conducted exclusively using 5 NVIDIA TITAN RTX graphics GPUs and 2 NVIDIA GeForce RTX 3090 GPUs, taking approximately 20 hours to complete. On average, 2 GPUs are allocated per experiment, amounting to a total of approximately 20400 GPU hours.
@@
Dataset Metrics Chat GPT GPT 4 Gemini LLaMA2 7B-chat LLaMA2 70B ChatGLM3 6B FinMA 7B FinGPT 7b-lora InternLM 7B Falcon 7B Mixtral 7B CFGPT sft-7B-Full
FPB F1 0.78* 0.78* 0.77 0.35 0.73 0.37 0.88 0.00 0.27 0.07 0.29 0.35*
Acc 0.78* 0.76* 0.77 0.29 0.72 0.38 0.88 0.00 0.20 0.05 0.37 0.26*
FiQA-SA F1 0.60 0.80 0.81 0.00 0.83 0.60 0.79 0.00 0.00 0.77 0.16 0.42*
[…]
Headlines AvgF1 0.77* 0.86* 0.78 0.60 0.63 0.60 0.97 0.60 0.60 0.45 0.60 0.61*
FOMC F1 0.64 0.71 0.53 0.53 0.49 0.47 0.49 0.00 0.00 0.30 0.37 0.16*
Acc 0.6 0.69 0.60 0.00 0.47 0.55 0.46 0.00 0.00 0.30 0.35 0.21*
FinArg-ACC MicroF1 0.50 0.60 0.31 0.51 0.58 0.50 0.27 0.00 0.47 0.23 0.39 0.05
FinArg-ARC MicroF1 0.39 0.40 0.60 0.28 0.36 0.30 0.08 0.00 0.37 0.32 0.57 0.05
MultiFin MicroF1 0.59 0.65 0.62 0.56 0.63 0.22 0.14 0.00 0.33 0.09 0.37 0.05
MA MicroF1 0.85 0.79 0.84 0.81 0.86 0.50 0.45 0.00 0.74 0.39 0.34 0.25
MLESG MicroF1 0.25 0.35 0.34 0.18 0.31 0.14 0.00 0.00 0.24 0.06 0.17 0.01
NER EntityF1 0.77* 0.83* 0.61 0.00 0.04 0.01 0.69 0.00 0.00 0.00 0.24 0.00
FINER-ORD EntityF1 0.28 0.77 0.14 0.00 0.07 0.00 0.00 0.00 0.00 0.00 0.05 0.00
@@
EDTSUM Rouge-1 0.17 0.20 0.39 0.11 0.25 0.14 0.13 0.00 0.12 0.15 0.12 0.01
BertScore 0.66 0.67 0.72 0.59 0.68 0.46 0.38 0.50 0.51 0.57 0.61 0.51
BartScore -3.64 -3.62 -3.87 -4.26 -3.81 -4.94 -5.71 -7.25 -4.51 -6.1 -4.47 -7.08
@@
BigData22 Acc 0.53 0.54 0.55 0.55 0.47 0.47 0.51 0.45 0.52 0.55 0.46 0.45
MCC -0.025 0.03 0.04 0.03 0.00 0.03 0.02 0.00 0.02 0.00 0.02 0.03
ACL18 Acc 0.50 0.52 0.52 0.51 0.51 0.49 0.51 0.49 0.52 0.51 0.49 0.48
MCC 0.005 0.02 0.04 0.01 0.01 0.00 0.03 0.00 0.02 0.00 0.00 -0.03
CIKM18 Acc 0.55 0.57 0.54 0.54 0.49 0.42 0.50 0.49 0.57 0.47 0.42 0.41
MCC 0.01 0.02 0.02 -0.01 -0.07 0.04 0.08 0.00 0.01 -0.06 -0.05 -0.07
German F1 0.20 0.55 0.52 0.52 0.17 0.52 0.17 0.52 0.44 0.23 0.53 0.53
MCC -0.10 -0.02 0.00 0.00 0.00 0.00 0.00 0.00 -0.01 -0.07 0.00 0.00
Australian F1 0.41 0.74 0.26 0.26 0.41 0.26 0.41 0.41 0.27 0.26 0.26 0.29
MCC 0.00 0.47 0.00 0.00 0.00 0.00 0.00 0.00 -0.02 0.00 0.00 -0.10
LendingClub F1 0.20 0.55 0.65 0.73 0.17 0.65 0.61 0.00 0.71 0.02 0.61 0.05
MCC -0.10 -0.02 0.19 0.04 0.00 0.19 0.00 0.00 0.18 -0.01 0.08 0.01
ccf F1 0.20 0.55 0.96 0.00 0.17 0.96 0.00 0.00 1.00 0.10 0.00 0.00
MCC -0.10 -0.02 -0.01 0.00 0.00 -0.01 0.00 0.00 0.00 0.00 0.00 0.00
ccfraud F1 0.20 0.55 0.90 0.88 0.17 0.90 0.01 0.00 0.77 0.62 0.48 0.03
MCC -0.10 -0.02 0.00 -0.05 0.00 -0.00 -0.06 0.00 -0.05 -0.02 0.16 0.01
polish F1 0.20 0.55 0.86 0.82 0.17 0.01 0.92 0.07 0.90 0.76 0.92 0.40
MCC -0.10 -0.02 0.14 0.08 0.00 0.01 -0.01 0.02 -0.01 0.05 0.00 -0.02
taiwan F1 0.20 0.55 0.95 0.75 0.17 0.95 0.95 0.82 0.48 0.00 0.95 0.70
MCC -0.10 -0.02 0.00 -0.07 0.00 0.00 0.00 0.01 0.04 0.00 0.00 0.00
portoseguro F1 0.20 0.55 0.95 0.01 0.17 0.97 0.04 0.00 0.95 0.95 0.72 0.00
MCC -0.10 -0.02 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.01 0.00
travelinsurance F1 0.20 0.55 0.00 0.00 0.17 0.00 0.00 0.98 0.79 0.77 0.00 0.03
MCC -0.10 -0.02 0.00 0.00 0.00 0.00 0.00 0.00 0.00 -0.03 0.00 0.01
Model CR (%)normal-$\uparrow$ SRnormal-$\uparrow$ DV (%)normal-$\downarrow$ AV (%)normal-$\downarrow$ MD (%)normal-$\downarrow$
Buy and Hold -4.83$\pm$18.9 0.0541$\pm$0.647 3.68$\pm$1.18 58.3$\pm$18.8 35.3$\pm$13
GPT-4 28.3$\pm$12.5 1.42$\pm$0.575 2.78$\pm$0.949 44.1$\pm$15 18.5 $\pm$ 6.92
ChatGPT 5.46$\pm$15.5 0.139$\pm$0.755 3.14$\pm$1.16 49.9$\pm$18.5 32.1$\pm$10.3
LLaMA2-70B 4.07$\pm$ 20.2 0.486$\pm$1.12 2.41$\pm$0.873 38.2$\pm$13.9 23.1$\pm$11.9
Gemini 21$\pm$21.6 0.861$\pm$0.832 2.5$\pm$ 1.23 39.7$\pm$19.6 22.5$\pm$7.9
Table 4: The average trading performance (95% Confidence Interval) comparison for different LLMs across 8 stocks. The results include large LLMs only ($\geq 70B$), as models with smaller contexts have difficulty understanding the instructions and producing a static strategy of holding.
4 Results
Table 3 and Table 4 shows the performance of 12 representative LLMs on all datasets in the FinBen.
4.1 Foundamental Tasks Analysis
@@
From Table 3, for fundamental tasks, we can see that GPT-4 stands out with the highest average performance, closely followed by ChatGPT, and Gemini. Among all open-sourced LLMs, the financial LLM FinMA-7B showcases superior performance on several classification tasks, such as FPB, even exceeding larger models like GPT-4. This is attributed to its tailored instruction tuning on the training datasets. For general-purpose LLMs, LLaMA2 70B leads in average performance, due to the large model size. Among models tailored for the Chinese language, ChatGLM2-6B outperforms InternLM 7B in average performance, indicating its effectiveness in handling financial tasks. However, CFGPT sft-7B-Full, fine-tuned on Chinese financial data, exhibits limited improvement on a few datasets and even declining performance on others like MultiFin compared to its base model InternLM 7B. This trend suggests a language-based discrepancy, highlighting that fine-tuning with Chinese data may adversely affect performance on English tasks, underscoring the complexities of cross-lingual adaptation in model training.
Notably, in quantification datasets such as Headlines, models like Gemini and other financially tuned LLMs, including FinMA-7B, perform on par with or even better than GPT-4. However, when tackling understanding tasks in datasets like FinQA and ConvFinQA, GPT-4 and ChatGPT significantly outperform others, highlighting the limited numerical reasoning capabilities of models like Gemini and LLaMA2-70B. Challenges persist in extraction datasets requiring complex information extraction and numeric labeling, such as FinRED, CD, FNXL, and FSRL, where all models, including GPT-4, fall short, indicating a need for further enhancement in these areas.
In conclusion, SOTA LLMs like GPT-4 exhibit strong performance across quantification tasks. However, there’s a clear gap in numerical reasoning and complex information extraction tasks, pinpointing the necessity for further development. Instruction tuning has shown to enhance performance significantly, suggesting a valuable approach for improving model capabilities in specialized financial tasks. The results highlight the complexity of cross-lingual model tuning and the importance of careful language consideration in enhancing LLMs’ effectiveness across diverse financial tasks.
4.2 Advanced Cognitive Engagement Tasks Analysis

### v2

@@ 2.3 Tasks
Text generation. This task assesses the models’ ability to produce coherent and informative text. Our focus is on text summarization, utilizing the ECTSUM (Mukherjee et al., 2022) dataset for summarizing earnings call transcripts. We also propose a novel dataset, EDTSUM, specifically designed for condensing financial news articles into concise summaries, constructed from original data in (Zhou et al., 2021). Evaluation employs ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), and BART Score (Yuan et al., 2021) to measure alignment, factual consistency, and information retention between machine-generated and expert summaries.
Forecasting. The forecasting task challenges models to adaptively predict future market and investor behaviors from emerging patterns. We focus on the1) Stock movement prediction task, forecasting stock directions as either positive or negative, based on historical prices and tweets, utilizing three datasets: BigData22 (Soun et al., 2022), ACL18 (Xu and Cohen, 2018) and CIKM18 (Wu et al., 2018).
@@
Risk management. It challenges LLMs to accurately identify, extract, and analyze relevant risk-related information, interpret numerical data, and understand complex relationships. We include 4 tasks: 1) Credit scoring classifies individuals as "good" or "bad" credit risks using historical customer data, employing datasets including: German (Hofmann, 1994), Australia (Quinlan, [n. d.]) and LendingClub (Feng et al., 2023). 2) Fraud detection involve categorizes transactions as "fraudulent" or "non-fraudulent", using two datasets: ccf (Feng et al., 2023) and ccFraud (Feng et al., 2023). 3) Financial distress identification aims to predict a company’s bankruptcy risk, using the polish (Feng et al., 2023) and taiwan dataset (Feng et al., 2023). Note that the dataset name describes only the region of the company, and the content within the datasets is in English. 4) Claim analysis anonymizes client data for privacy, labeling a "target" to indicate claim status, using two datasets: PortoSeguro (Feng et al., 2023) and travelinsurance (Feng et al., 2023). It is noticed that the dataset name such as German and taiwan, only indicates customer sources and all content is in English. F1 score and Matthews correlation coefficient (MCC) (Chicco and Jurman, 2020) are used for evaluating these tasks.
@@
Decision-making. Strategic decision-making (Punt, 2017) evaluates the model’s proficiency in synthesizing diverse information to formulate and implement trading strategies, a challenge even for experts. We innovatively introduce the SOTA financial LLM agent FinMem (Yu et al., 2023) to evaluate LLMs on the stock trading task. We construct the novel FinTrade dataset, containing 10 stocks, simulating real-world trading through historical prices, news, and sentiment analysis. Performance is measured by Cumulative Return (CR) (Ariel, 1987), Sharpe Ratio (SR) (Sharpe, 1998), Daily (DV) and Annualized volatility (AV) (Zhou et al., 2023), and Maximum Drawdown (MD) (Magdon-Ismail and Atiya, 2004), offering a comprehensive assessment of profitability, risk management, and decision-making prowess.
@@ 3 Evaluation
We evaluate the zero-shot (from our evaluation) and few-shots (results from previous papers) performance of 15 representative general LLMs and financial LLMs on the FinBen benchmark, including: 1) ChatGPT: A LLM developed by OpenAI. 2) GPT-4 (OpenAI, 2023b): The SOTA commercialized LLMs proposed by OpenAI. 3) Gemini Pro Team et al. (2023): A multimodal AI LLM with 50T parameters, released by Google. 4) LLaMA2-7/70B-chat Touvron et al. (2023b): An open-sourced instruction-following LLM with 7B and 70B parameters developed by MetaAI. 5) LLaMA3-8B [^5] https://llama.meta.com/llama3/: An open-sourced LLMs developed by MetaAI, using more training data than LLaMA2. 6) ChatGLM3-6B Du et al. (2022): A conversational LLM with 6B parameters, jointly released by Zhipu AI and Tsinghua KEG. 7) Baichuan2-6B Baichuan (2023): An open-source LLM with 6B parameters, launched by Baichuan Intelligent Technology. 8) InternLM-7B Team (2023): An open-sourced 7B parameter base model tailored for practical scenarios, proposed by SenseTime. 9) Falcon-7B Almazrouei et al. (2023): A 7B parameter causal decoder-only LLM model trained on 1500B tokens of RefinedWeb enhanced with curated corpora. 10) Mixtral 8$\times$7B Jiang et al. (2024): A LLM with the Sparse Mixture of Experts (SMoE) architecture. 11) Code LLaMA-7B Roziere et al. (2023): An open-source LLM model for generating programming code, launched by Meta AI with 7B parameters. 12) FinGPT Yang et al. (2023a): A 7B instruction finetuned financial LLM based on LLaMA 7B Touvron et al. (2023a) with sentiment analysis tasks. 13) FinMA-7B Xie et al. (2023b): A 7B instruction finetuned financial LLM based on LLaMA 7B with multiple NLP and forecasting tasks. 14) DISC-FinLLM Chen et al. (2023b): An open-sourced financial LLM, fine-tuned from Baichuan-13B-Chat Baichuan (2023). 15) CFGPT Li et al. (2023a): An open-source LLM, specifically designed for the financial sector and trained on Chinese financial datasets, which comprises 7B parameters.
Experimental Settings We set the maximum generation tokens for LLMs to 1024 and the batch size to 20,000 for all experiments. These experiments are exclusively conducted on 16 NVIDIA A100 80G GPUs, taking approximately 600 hours to complete. Including the GPT-4 API costs, the total expenditure amounts to approximately \$51,000.
4 Results
Table 3 and Table 4 shows the performance of 12 representative LLMs on all datasets in the FinBen. We also report results of non-LLM methods (traditional methods) in Appendix G.
4.1 Information Extraction & Textual Analysis Results
As shown in Table 3, for IE tasks, GPT-4 demonstrates superior performance in named entity recognition tasks, including NER, FINER-ORD, and FinRED. InternLM 7B achieves the best results in causal classification (SC). However, for more complex information extraction tasks, such as causal detection (CD) and numerical understanding (FNXL and FSRL), even GPT-4’s performance is limited, with Gemini showing only slightly better results, still falling short of expectations. Additionally, while financial domain-specific LLMs developed by instruction tuning such as FinMA 7B exhibit improvements over general domain LLMs such as LLaMA2 7B-chat, they continue to struggle with both named entity recognition and complex extraction tasks. These findings highlight significant opportunities for advancement in financial causal detection and numerical understanding for LLMs.
@@ 4.1 Information Extraction & Textual Analysis Results
Dataset Metrics Chat GPT GPT 4 Gemini LLaMA2 7B-chat LLaMA2 70B LLaMA3 8B FinMA 7B FinGPT 7b-lora InternLM 7B Falcon 7B Mixtral 7B CFGPT sft-7B-Full
NER EntityF1 0.77* 0.83* 0.61 0.18 0.04 0.08 0.69 0.00 0.00 0.00 0.24 0.00
FINER-ORD EntityF1 0.28 0.77 0.14 0.02 0.07 0.00 0.00 0.00 0.00 0.00 0.05 0.00
@@
FPB F1 0.78* 0.78* 0.77 0.39 0.73 0.52 0.88 0.00 0.69 0.07 0.29 0.35*
Acc 0.78* 0.76* 0.77 0.41 0.72 0.52 0.88 0.00 0.69 0.05 0.37 0.26*
FiQA-SA F1 0.60 0.80 0.81 0.76 0.83 0.70 0.79 0.00 0.81 0.77 0.16 0.42*
[…]
Headlines AvgF1 0.77* 0.86* 0.78 0.72 0.63 0.60 0.97 0.60 0.60 0.45 0.60 0.61*
FOMC F1 0.64 0.71 0.40 0.35 0.49 0.40 0.49 0.00 0.36 0.30 0.37 0.16*
Acc 0.6 0.69 0.60 0.49 0.47 0.41 0.46 0.00 0.35 0.30 0.35 0.21*
FinArg-ACC MicroF1 0.50 0.60 0.31 0.46 0.58 0.51 0.27 0.00 0.39 0.23 0.39 0.05
FinArg-ARC MicroF1 0.39 0.40 0.60 0.27 0.36 0.28 0.08 0.00 0.33 0.32 0.57 0.05
MultiFin MicroF1 0.59 0.65 0.62 0.20 0.63 0.39 0.14 0.00 0.34 0.09 0.37 0.05
MA MicroF1 0.85 0.79 0.84 0.70 0.86 0.34 0.45 0.00 0.78 0.39 0.34 0.25
MLESG MicroF1 0.25 0.35 0.34 0.03 0.31 0.12 0.00 0.00 0.14 0.06 0.17 0.01
TATQA EmAcc 0.00* 0.13* 0.18 0.03 0.01 0.01 0.00 0.00 0.00 0.00 0.01 0.00
Regulations Rouge-1 0.12 0.11 - 0.24 - 0.10 0.12 0.01 0.04 0.03 - 0.14
BertScore 0.64 0.62 - 0.65 - 0.60 0.59 0.40 0.57 0.14 - 0.57
EDTSUM Rouge-1 0.17 0.20 0.39 0.17 0.25 0.14 0.13 0.00 0.13 0.15 0.12 0.01
BertScore 0.66 0.67 0.72 0.62 0.68 0.60 0.38 0.52 0.48 0.57 0.61 0.51
BartScore -3.64 -3.62 -3.87 -3.99 -3.81 -4.94 -5.71 -7.23 -4.60 -6.1 -4.47 -7.08
@@
BigData22 Acc 0.53 0.54 0.55 0.54 0.47 0.55 0.51 0.45 0.56 0.55 0.46 0.45
MCC -0.025 0.03 0.04 0.05 0.00 0.02 0.02 0.00 0.08 0.00 0.02 0.03
ACL18 Acc 0.50 0.52 0.52 0.51 0.51 0.52 0.51 0.49 0.51 0.51 0.49 0.48
MCC 0.005 0.02 0.04 0.01 0.01 0.02 0.03 0.00 0.02 0.00 0.00 -0.03
CIKM18 Acc 0.55 0.57 0.54 0.55 0.49 0.57 0.50 0.42 0.57 0.47 0.42 0.41
MCC 0.01 0.02 0.02 -0.03 -0.07 0.03 0.08 0.00 -0.03 -0.06 -0.05 -0.07
German F1 0.20 0.55 0.52 0.57 0.17 0.56 0.17 0.52 0.41 0.23 0.53 0.53
MCC -0.10 -0.02 0.00 0.03 0.00 0.05 0.00 0.00 -0.30 -0.07 0.00 0.00
Australian F1 0.41 0.74 0.26 0.26 0.41 0.26 0.41 0.38 0.34 0.26 0.26 0.29
MCC 0.00 0.47 0.00 0.00 0.00 0.00 0.00 0.11 0.13 0.00 0.00 -0.10
LendingClub F1 0.20 0.55 0.65 0.72 0.17 0.10 0.61 0.00 0.59 0.02 0.61 0.05
MCC -0.10 -0.02 0.19 0.00 0.00 -0.15 0.00 0.00 0.15 -0.01 0.08 0.01
ccf F1 0.20 0.55 0.96 0.00 0.17 0.01 0.00 1.00 1.00 0.10 0.00 0.00
MCC -0.10 -0.02 -0.01 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
ccfraud F1 0.20 0.55 0.90 0.25 0.17 0.36 0.01 0.00 0.57 0.62 0.48 0.03
MCC -0.10 -0.02 0.00 -0.16 0.00 -0.03 -0.06 0.00 -0.13 -0.02 0.16 0.01
polish F1 0.20 0.55 0.86 0.92 0.17 0.83 0.92 0.30 0.92 0.76 0.92 0.40
MCC -0.10 -0.02 0.14 0.00 0.00 -0.06 -0.01 0.00 0.07 0.05 0.00 -0.02
taiwan F1 0.20 0.55 0.95 0.95 0.17 0.26 0.95 0.60 0.95 0.00 0.95 0.70
MCC -0.10 -0.02 0.00 -0.01 0.00 -0.07 0.00 -0.02 -0.01 0.00 0.00 0.00
portoseguro F1 0.20 0.55 0.95 0.01 0.17 0.94 0.04 0.96 0.96 0.95 0.72 0.00
MCC -0.10 -0.02 0.00 -0.05 0.00 -0.01 0.01 0.00 0.00 0.00 0.01 0.00
travelinsurance F1 0.20 0.55 0.00 0.00 0.17 0.00 0.00 0.98 0.89 0.77 0.00 0.03
MCC -0.10 -0.02 0.00 0.00 0.00 0.00 0.00 0.00 0.12 -0.03 0.00 0.01
Regarding TA tasks, instruction fine-tuned models like FinMA 7B exhibit the best performance in sentiment analysis tasks, including FPB, FiQA-SA, and Headlines. However, the generalization ability of FinMA 7B is limited due to the diversity of TA tasks in the financial domain. It performs even worse than general domain LLMs such as LLaMA2-7B-chat on other TA tasks, where GPT-4, Gemini, and LLaMA2 70B show superior results. This underscores the limitations of instruction fine-tuned models, which may also be constrained by the parameter size of their backbone models.
@@
Models tailored for the Chinese language, such as CFGPT sft-7B-Full, which is fine-tuned on Chinese financial data, exhibit limited improvement on some datasets and even a decline in performance on others like MultiFin compared to its base model InternLM 7B. This trend suggests a language-based discrepancy, indicating that fine-tuning with Chinese data may adversely affect performance on English tasks. These findings underscore the complexities of cross-lingual adaptation in model training, highlighting the challenges in achieving consistent performance across different languages.
4.2 Question Answering & Text Generation Results
In the QA tasks, closed-source commercial LLMs like GPT-4 and Gemini continue to lead across all datasets. While FinMA 7B shows improvement over its backbone models, it remains limited by model size and exhibits bottlenecks in numeric reasoning ability. For the regulations dataset, which is the first intersection dataset requiring both financial and legal knowledge, GPT-4 demonstrates its broad knowledge coverage effectively.